# From Imitation to Refinement – Residual RL for Precise Assembly

### v3

@@ II Problem Setup and Approach Overview
TABLE I: Top BC-trained MLPs without chunking (MLP-S) cannot perform any of the tasks, and Diffusion Policies (DP) generally outperform MLPs with chunking (MLP-C). Bottom Training our proposed residual policies with RL on top of frozen DP performs the best among the evaluated fine-tuning techniques. The best checkpoint is chosen for all methods, and we average over 1024 rollouts to get the success rate.
III Methods
III-A Imitation Learning
For each task, we collect a dataset of 50 demonstrations in simulation by teleoperating the robot, $\mathcal{D}_{\text{sim}}:\ \{\tau_{1},...,\tau_{N}\}$ to train a “base” policy that serves as the starting point for RL training. Each trajectory, $\tau$, contains system states $s_{t}$, and robot actions $a_{t}$, i.e., $\tau_{i}=\{(s_{t},a_{1}),...,(s_{T},a_{T})\}$, with $T$ being the trajectory length.
$\mathcal{D}_{\text{sim}}$ is used to train base policy $\pi_{\text{base}}$ with Behavior Cloning (BC), i.e., $\max_{\pi}\ \mathbb{E}_{(a_{t},s_{t})\sim\mathcal{D}_{\text{sim}}}\left[\log\pi(a_{t}|s_{t})\right]$. We use a Diffusion Policy [18] as the base model, which has shown strong empirical performance in handling difficult manipulation tasks with relatively small datasets [21]. Consistent with recent advancements [18, 6], our policy framework enhances its performance by predicting multiple future actions in chunks instead of individual actions at each timestep. We denote the length of future action sequences predicted by the policy as $T_{a}$, the output as $\mathbf{a_{t}}=[a_{t},...,a_{t+T_{a}}]$. When predicting an action chunk $\mathbf{a_{t}}$ of length $T_{a}$, we only execute a subset $[a_{t},...,a_{t+T_{\text{exec}}}]$, with execution horizon $T_{\text{exec}}\leq T_{a}$. We typically use $T_{a}=32$ and $T_{\text{exec}}=8$.
III-B Online Reinforcement Learning with Residual Policies
Given the initial base policy $\pi_{\text{base}}$, we aim to improve it using RL. However, directly fine-tuning diffusion models with RL is an active area of research [55, 56, 46, 57, 47], made difficult because of the multi-step inference process and unavailability of the policy action log-probabilities. Another category of methods upweight high-quality model outputs via importance sampling, return conditioning, or augmenting the original de-noising objective with a loss term for maximizing a Q-function [58, 59, 60, 17, 61]. However, these methods mainly enable better extraction of high-quality behavior in the data, while we are more concerned with learning new, corrective behaviors. Action chunks can also make optimization with policy gradients more difficult. This is partly due to chunking increasing the effective action space (e.g., chunks of 8 actions increase the dimension 8$\times$), which we find to increase RL training instability (see Sec. IV-A). This means fine-tuning other popular BC architectures, like the Action-Chunked Transformer (ACT) [6], also brings technical challenges [15]. Even with successful RL fine-tuning, the improved policy will still be a trajectory planner lacking in closed-loop reactive control (Fig. 4). Finally, recent work has shown that fine-tuning large pre-trained models can lead to forgetting of capabilities if the agent does not visit states seen during pre-training frequently enough [62].
We side-step these complications by training a residual Gaussian Multi-Layer Perceptron (MLP) [63] policy $\pi_{\text{res}}$. This policy takes as input both the system state and the action predicted by the diffusion policy $\pi_{\text{base}}$ and produces an “action correction” that modifies the action as $a_{t}=a_{t}^{\text{base}}+\alpha\cdot a_{t}^{\text{res}}$, with $\alpha\leq 1$ being a coefficient controlling what scale the residual policy operates on. We find that the policy is not sensitive to $\alpha$ in isolation, but rather the resulting exploration noise, $\alpha\mathcal{N}(0,\sigma^{\text{init std}}I)$, see App. I-C for more analysis. We denote the resulting combined policy $\boldsymbol{\pi}$. This decoupling has several advantages. First, we can regard the diffusion model as a black box and not change its parameters [48]. This decoupling also means that ACT (see App. I-B) and other BC methods can serve as the base model without changing the method – an advantage of residual policy learning emphasized by prior robotics work [48, 64]. This also allows different prediction horizons $T_{a}$ for the two policies. This flexibility in action horizon is helpful as most RL algorithms optimize single-action policies and allow the residual to be fully closed-loop. Finally, it removes the need to explicitly regularize the fine-tuned policy to stay near the pre-trained policy, which is often necessary to achieve stable optimization [65, 66].
Given the above, we train the $\pi_{\text{res}}$ using standard PPO [26] with the alternation that we augment the state space with the base action. The base model observes the current state $s_{t}$ and outputs a chunk of actions $\mathbf{a_{t}}{}$. For each action in the chunk, we concatenate it with the current observation $s_{t+i}^{\text{res}}=[s_{t+i},a^{\text{base}}_{t+i}]$, and predict the correction $a_{t+i}^{\text{res}}\sim\pi^{\text{res}}(\cdot|s_{t+i}^{\text{res}})$ for $i=0,...,T_{a}-1$. Besides the technical convenience, one can look at residual policy as a way to incorporate an inductive bias that the policy should mostly learn locally corrective actions [64]. This is motivated by our qualitative observations that most of the base BC policy failures are due to slight imprecision near “bottleneck” transitions when performing skills like grasping and insertion (e.g., LABEL:fig:overview (Left)).
III-C Sim-to-real pipeline
Ultimately, we are interested in what capabilities our method can enable on real robots. To that end, we devise a simple sim-to-real pipeline to transfer state-based policies trained in simulation to RGB-based policies deployable on physical hardware. We present the general pipeline in Fig. 3. Once the RL training has converged, the combined policy has attained closed-loop correction behavior. We aim to distill this enhanced performance from the state-based policy into a vision-based policy that operates from RGB images. Following the established teacher-student distillation paradigm [31, 67, 68], we generate a dataset of successful trajectories, $\mathcal{D}_{\text{synth}}$, where the environment’s observations, $o_{t}$, replace the states, i.e., $\mathcal{D}_{\text{synth}}=\{\tau_{\text{synth},1},...,\tau_{\text{synth},N}\}$ and $\tau_{\text{synth},i}=\{(o_{1},a_{1}),...,(o_{T},a_{T})\}$. For real-world transfer, we enhance the synthetic dataset $\mathcal{D}_{\text{synth}}$ by re-rendering its trajectories in Isaac Sim [69]. This process improves image quality and introduces variability in environmental conditions such as object and table colors, textures, lighting, and camera perspectives, many of which cannot easily be done with standard image augmentation techniques. We denote this refined dataset as $\mathcal{D}_{\text{synth-render}}$. To ease the difficulty of zero-shot sim-to-real with RGB images, we opt for a co-training approach, integrating the synthetic dataset with a small set of real-world task demonstrations, $\mathcal{D}_{\text{real}}$, which similarly comprises only environmental observations without ground truth poses. This combined dataset, $\mathcal{D}_{\text{real}}\cup\mathcal{D}_{\text{synth-render}}$, is used to train the final student policy with BC.
IV Experiments and Results
@@ 40 Real + 350 Sim 10/10 10/10 9/10 10/10 6/10 7/10 5/10 6/10 5/10 6/10
TABLE II: We investigate the effect of combining real-world demonstrations with simulation trajectories from our RL-trained residual policies. Co-training with real and synthetic data improves motion quality and success rate on the one_leg task.
Figure 4: Perturbations highlight ResiP’s resilience: 12% drop vs. 19-26% for chunked methods. See App. I-A for more detailed discussion.
First, we investigate the requirement of complex BC methods and the impact of our residual reinforcement learning approach on improving the success rates of policies trained with imitation learning in Sec. IV-A. Next, we study the performance of downstream distillation from synthetic RL data, examining the relationship between the quantity and quality of the data and the resulting performance of the distilled vision-based policies in Sec. IV-B. Finally, we apply our approach to real robot hardware, enabling precise assembly tasks on a physical robot directly from vision input in Sec. IV-C.
IV-A Improving Imitation Learning with Residual Learning
Why Action Chunking and Diffusion Policies?
Another interesting finding is that the DP is significantly more robust to noise, an important property for an adequate base policy for residual fine-tuning. In contrast, we observed performance collapse when using MLP-C as the base, further reinforcing the utility of using powerful generative models for BC pre-training. See more analysis in App. I-B.
Online Policy Fine-Tuning Comparison
Next, we investigate the effectiveness of our residual RL approach, ResiP, in improving the performance of policies trained with BC. We train residual policies using PPO [26] on top of the diffusion policy and compare it to two baselines – (1) directly fine-tune the MLP-C with PPO while regarding every action chunk as one concatenated action and (2) a chunked version of the IDQL algorithm [58]. In the IDQL algorithm, multiple action chunks are sampled from the diffusion policy, and one chooses the action chunk to execute based on its Q-value. We have found that learning a good Q-value estimator from offline data alone, as proposed in IDQL, is challenging with only 50 demonstrations. Therefore, we learn it in on-policy matter similar to [70].
The results in Tab. I show that our residual RL approach significantly improves success rates over BC and outperforms alternative RL fine-tuning methods. For tasks with lower initial randomization, such as the one_leg task, the residual policy increases the success rate from 54% to 98%. Even more drastically, for the peg-in-hole task, ResiP improved the success rate from 5% to 99%. However, we observe performance saturates at lower success rates (e.g., $\sim$70%) for tasks with higher part randomization. We hypothesize that the base model’s performance limits the residual policy as it is designed to act locally, constrained by the base.
Fig. 5 qualitatively shows the most common failure mode of the BC policies. In LABEL:fig:overview (Left), a specific failure mode and how the RL policy overcomes it is further exemplified. For instance, a typical BC policy error is to push down before the leg is aligned with the hole. This often results in a shift of the object in the grasp, which causes the policy to diverge due to the out-of-distribution grasp pose. The residual policy reliably corrects these errors by performing small sideways translations while canceling premature downward motions. It typically only allows the leg to be inserted once it is appropriately above the hole. We also find the residual policy is better at performing initial grasps that allow accurate downstream alignment between the grasped object and the receptacle. On the other hand, the BC policies more often grasp the object at angles that make insertion more difficult. In Fig. 6, we qualitatively analyze this and find that the residual makes the precise insertion part of the task significantly more reliable.
Figure 5: Examples of common failures in tasks that require precise part alignments. All of these failures typically only require small corrections to ameliorate, making the residual framework well-suited to improve on the nominal policy.
Surprisingly, the residual RL procedure also leads to the emergence of qualitatively different behaviors. We refer to the website’s uncut videos of rollouts for the pre-trained and fine-tuned policies. In these videos, we observe that the fine-tuned policy has learned new grasps of the table foot that were not demonstrated, as well as more persistent retrying behavior and recovery.
In addition to improved success rates, we observed different training dynamics across each method. First, direct MLP-C fine-tuning proved unstable and required KL-regularization to avoid collapse. Second, the trained DP produced actions with low variance, even with $\eta=1$ in the DDIM sampler [71], inhibiting Q-learning and constraining the potential for policy improvement. Finally, residual policy training was very stable, likely because it is constrained to operate on a local scale, which prevents large deviations that can make RL unstable and has a small parameter count compared to the base policy [72, 73].
What drives residual RL performance?
To further understand what drives the residual RL fine-tuning approach’s performance, we try to decompose it into an on-policy component and a closed-loop, reactive component.
The impact of on-policy learning is probed by comparing learning the task by scaling the purely offline data versus learning online with DAgger [74]. The performance implications of distillation are further addressed in Sec. IV-B. We find that when applying DAgger for the same action-chunked diffusion policy, we can improve performance significantly over the purely offline case, as shown in LABEL:fig:overview (right) and Fig. 7. However, a gap remains between the residual RL and DAgger policy of 9% and 23% for the one_leg and round_table tasks, respectively.
On the other hand, we also find the residual policy’s closed-loop, single-step nature helpful. First, when trying to learn a residual policy for whole chunks, the resulting policy learns significantly more slowly, and the performance saturates at a lower level than that of the full closed-loop policy. See App. I-A for details. Furthermore, as shown in Fig. 4, the effect of fully closed-loop control becomes clear when we introduce randomly sampled force perturbations to the episode rollouts. Performance for the ResiP policy drops 12 percentage points, compared to 19-26 percentage points for alternative chunked methods (BC, DAgger, chunked residual).
Figure 7: When scaling up the offline dataset for the harder round_table task, the same trends emerge as for one_leg (seen in LABEL:fig:overview (right)), but with saturation happening at a lower success rate of 56%. Again, correcting for distribution shifts with DAgger improves performance but saturates at 71%.
IV-B Distillation Performance from Synthetic RL Data
Next, we study how synthetic RL data quantity and quality impact the performance of distilled vision-based policies. First, we find that distilling trajectories from the RL agent performs better than training directly on the 50 demonstrations. A vision-based policy distilled from the RL-trained teacher reached 73% on one_leg, outperforming the 50% achieved by training the vision policy directly on human demos, see Fig. 8. However, we also observe a performance gap between the RL-trained teacher (98%) and distilled vision-based student policy (73%). We consider whether this gap may be caused by training the student to operate on images. Upon examination, we find that distilling the same number of teacher rollouts into an image-based student and a state-based student results in comparable performance. This leads us to conclude that the change in modality is not the primary source of the performance gap.
Figure 8: Comparison of distilled performance from BC and RL-based teacher. On the left, BC policies trained on the same set of demonstrations achieve comparable performance for state and image observations. On the right, the vision-based student policy has a performance gap similar to the state-based teacher for state-to-state distillation.
Therefore, we also examine the impact of the distillation dataset size. Here, we scale up the number of state-based rollouts from the trained RL policy and distill these to a state-based student. In LABEL:fig:overview (Right), we see that performance increases with more data but still does not reach the performance achieved by the teacher policy, with a gap of 20 percentage points between the best student policy (80% success rate) and the teacher policy (98% success rate) at 100k trajectories (with only a minor improvement from 78% at 10k trajectories). The same trend is evident in Fig. 7 (though the saturation occurs earlier). Nevertheless, the improved performance obtained by training with more data highlights the advantage of using simulation for obtaining large-scale synthetic datasets.
DAgger [1] demonstrates significantly greater sample and computational efficiency than pure BC. As shown in LABEL:fig:overview (right), DAgger achieves superior performance to BC trained on 50 human demos in just $\sim$10k gradient steps (with 800 rollouts) and surpasses our best BC distillation results at $\sim$20k steps (with $\sim$1.5k demonstrations). These results underscore the advantages of interactive learning over passive offline learning and point towards effective online distillation techniques to help close the gap between the state-based RL-trained experts and the distilled real-world policies deployed on real hardware (see Sec. IV-C).
IV-C Real-World Application
Figure 9: Examples of photo-realistic rendering and domain randomization for sim-to-real transfer. In the first row, we show our nominal experiment setting of rendering parts in the original white color while varying the position, brightness, and hue of the lighting in the scene. In the bottom row, we also introduce variations in the part colors to see how that can help the final policy adapt to unseen part appearances without collecting any additional data.
@@ IV-C Real-World Application
Finally, we evaluate the real-world performance of a sim-to-real policy trained on a mixture of a few (10/40) real-world demonstrations and simulation data generated by the trained residual RL policy. We compare the co-trained policy to a baseline model trained only on real-world demonstrations. We compare the success rates achieved by each policy on two sets of 10 trials for the one_leg task. In the first set, we randomize part poses, while in the second set, we randomize obstacle poses (i.e., insertion location in the workspace).
The results, presented in Tab. II, show that incorporating simulation data improves real-world performance (e.g., increasing task completion rate from 20-30% to 50-60%). Qualitatively, the sim-to-real policy exhibits smoother behavior and makes fewer erratic movements that might exceed the robot’s physical limits. Fig. 10 shows examples of successful and unsuccessful task attempts.
To further probe the sim-to-real transfer and potential for acquiring new capabilities with the help of the simulator, we create a task variation where the part colors are changed from black to white. When rolling out the policy trained on real demos of white parts, the robot exhibited erratic behavior that caused the hardware to reach velocity limits on every trial we ran, as shown in Fig. 11 (A). When including synthetic data rendered with parts in black, the resulting policy can perform the task again (see Fig. 11). The resulting performance was still inferior to the performance on white parts, which motivates further work on closing the sim-to-real gap.

### v4

@@ III-A Tasks and Environment
TABLE I: Success rates (percentage of successful completions over 1024 evaluation episodes using each method’s best-performing checkpoint) for different policy architectures across our task suite. Top: Baseline BC methods, where Diffusion Policies (DP) generally outperform MLPs with chunking (MLP-C), while MLPs without chunking (MLP-S) are unable to solve all tasks except peg-in-hole. Bottom: RL-based methods, where our proposed residual policy (ResiP) provides significant improvements in success rates over the base BC policy, improvements not seen from the alternative RL methods.
III-B System Configuration
The policy operates at 10Hz on a 7 Degrees-of-Freedom (DoF) Franka Emika Panda robot arm for all tasks but biman-insert, which operates at 50Hz on two Franka Panda arms (i.e., 14 DoF). The action space consists of the desired end-effector pose $\mathbf{T}{}^{\text{des}}\in$ $\text{SE}(3)$ (i.e., both position and orientation) and a binary gripper command for opening/closing the parallel-jaw gripper. These desired end-effector poses are converted to joint position targets using differential inverse kinematics [63], then tracked using a low-level PD controller running at 1KHz with manually specified stiffness and dampening parameters lightly tuned to balance compliance with accurate trajectory tracking.
III-C Evaluation protocol
III-C1 Primary Evaluation
We evaluate all methods by executing the policy from an initial state sampled from the same initial state distribution used for collecting demonstrations. We follow the default randomization protocol used in FurnitureBench [5]: at the beginning of each episode, objects are randomly offset from their nominal positions by $\Delta x,\Delta y\in[-1.5,1.5]$ cm in the horizontal plane (with fixed height $z$) for low randomization settings and by $\Delta x,\Delta y\in[-5,5]$ cm for medium randomization. For the U-shaped fixture used to stabilize parts during assembly (the three-sided white “wall” surrounding the parts in the left-most three columns in Fig. 4), we apply additional offsets of $\Delta x,\Delta y\in[-2,2]$ cm and $\Delta x,\Delta y\in[-4,4]$ cm from its nominal position for low and medium settings respectively.
We define task success as achieving the target geometric alignment between pairs of parts, as specified by the task reward function in Sec. II-A. For each method, we report the success rate calculated over 1024 evaluation episodes using the best-performing checkpoint, as our goal is to demonstrate the potential capabilities of each architectural approach rather than average performance across training runs.
III-C2 Robustness to Dynamic Disturbances
To further evaluate robustness to dynamic disturbances, we also perform a version of the above evaluation that includes random force perturbations during policy execution. At each timestep, we randomly sample 1% of the objects in the environment and apply a perturbation corresponding to the same randomization described above for initialization. These perturbations simulate unexpected contact forces and dynamic disturbances not seen during training. We evaluate each method’s performance degradation under these conditions by comparing success rates with/without perturbations across 1024 rollouts.
III-C3 Real-World Evaluation Protocol
For sim-to-real transfer, we use IsaacSim [64] to render photorealistic trajectories of our simulation data, as it provides better rendering capabilities than IsaacGym [61]. We evaluate policies across a grid of initial object and obstacle poses for real-world experiments that match our low randomization simulation setting ($\Delta x,\Delta y\in[-1.5,1.5]$ cm). We perform 10 trials for each method and ensure that each method is tested on the same initial object states. Success criteria remain consistent with our simulation experiments: achieving target geometric alignment between assembly parts.
III-D Baselines and Ablations
We evaluate our method against several baselines to analyze the importance of action chunking, policy architecture, and closed-loop control learned with Reinforcement Learning (RL). Implementation details and hyperparameters are provided in App. A.
III-D1 Behavior Cloning Baselines
We first evaluate the impact of action chunking by comparing a standard Behavior Cloning (BC) approach using an MLP (MLP-S) with a version that predicts action chunks (MLP-C). We find the Diffusion Policy architecture [19] provides the strongest BC performance and use it as both our primary baseline and the foundation for ResiP.
III-D2 Distribution Shift Analysis
To assess the impact of distribution shifts, we implement DP-DAgger [1], which iteratively queries an expert policy to gather corrective demonstrations in states visited by the learned policy. Our experiments use ResiP as the expert. The DAgger policy is trained with the same BC loss and architecture as the nominal DP policy.
III-D3 Reinforcement Learning Comparisons
Building on our BC policies, we compare two approaches for RL fine-tuning. First, we evaluate PPO fine-tuning of our chunked MLP policy (PPO-C) [52], treating each action chunk as a single concatenated action. For our diffusion-based policy, we implement IDQL [65], where multiple action chunks are sampled from the diffusion policy and selected based on learned Q-values using the on-policy method of [66].
Figure 5: Scaling behavior comparison on the harder round_table task mirrors the trends observed for one_leg (LABEL:fig:overview (right)), but with lower overall performance. Just as in one_leg, increasing the number of demonstrations for BC training leads to diminishing returns, saturating at 56% success rate. While DAgger’s online data collection performs better, it also plateaus at 71%, well below ResiP’s 94% success rate. This consistent pattern across tasks, with lower saturation points for the more complex tasks, highlights the limitations of offline approaches and the benefits of learning closed-loop control online.
III-D4 Closed-Loop Control Ablation
To study the importance of per-timestep corrections, we implement ResiP-C, a variant of our method that predicts residual corrections at the same frequency as the base policy’s action chunks. While the standard ResiP observes the current state and predicts a correction for each timestep, ResiP-C observes the current state and all actions in the predicted chunk and predicts a correction of all actions in the chunk. ResiP-C uses the same online PPO training procedure as ResiP. This chunked correction makes the learning problem more challenging as the residual policy predicts corrections for future timesteps without access to the intermediate states. See Sec. J-A for details.
III-D5 Real-World Baselines
We compare the real-world performance of two policies: (1) Real-Only: Diffusion policies trained exclusively on real-world demonstrations $\mathcal{D}_{\text{real}}$, using either 10 or 40 demonstrations, e.g., denoted 10 Real-Only. (2) Real+Sim: Following our sim-to-real approach described in Sec. II-D, we combine the same real-world demonstrations with our synthetic rendered dataset, e.g., denoted 10 Real+Sim.
IV Experimental Results
Our experimental evaluation focuses on three key aspects. First, we analyze how augmenting chunked Behavior Cloning (BC) policy with closed-loop residual Reinforcement Learning (RL) enables reliable execution of precision-critical manipulation tasks (Sec. IV-A). Second, through ablation studies, we identify design choices crucial for ResiP’s performance gains (Sec. IV-B). Finally, we evaluate our method on physical robot hardware (Sec. IV-C), demonstrating improved real-world performance through teacher-student distillation while analyzing distillation bottlenecks.
IV-A Augmenting Trajectory Planning with Reactive Control
In simulation experiments, we evaluate the fundamental limitations of state-of-the-art BC methods on precision-critical tasks. Using a suite of 6 assembly tasks and 50 demonstrations per task, we find that the Diffusion Policy (DP) architecture struggles with high-precision requirements. Using the round_table task as an example, DP achieves a 12% success rate with 50 demonstrations. Increasing the number of demonstrations naturally improves performance. However, as shown in Fig. 5, scaling to 10,000 demonstrations only improves performance to 56% on the round_table task. Another potential solution is addressing the distribution shift through online data collection. While our DAgger implementation shows improved performance over pure BC, it plateaus at 71% success on the round_table task, still falling significantly short of expert-level performance (similar scaling behavior is observed for the one_leg task, see LABEL:fig:overview (Right)).
Our residual learning approach, ResiP, addresses these limitations with closed-loop control learned with RL. Starting from the same 50 demonstrations, ResiP achieves 94% on round_table (up from 12%) and 98% success on one_leg (up from 54%), showing significant gains over the alternatives. The most dramatic improvement is seen in the peg-in-hole task, where success rates increase from 5% to 99%, highlighting the particular effectiveness of our method when precise local corrections are the primary challenge. Tab. I shows comparisons across our task suite, demonstrating ResiP’s consistent advantages over both BC baselines and alternative RL methods.
Figure 6: Examples of the common failures of our DP policies in tasks requiring precise part alignments. These failures can be mitigated by small corrections, making our residual framework, ResiP, well-suited to improve the reliability of the nominal DP policy.
IV-A1 Analyzing Failure Modes
To understand where the large performance improvements stem from, we analyzed failure modes of the DP policies, shown in Fig. 6 and LABEL:fig:overview (Left). In the low randomization setting, we observe that DP’s failures primarily arise from small imprecisions: a common error is pushing the leg down before achieving perfect alignment with the hole. Consequently, the object’s pose shifts slightly in the gripper, causing an out-of-distribution grasp pose. The residual policy reliably corrects these errors through minimal adjustments: performing small sideways translations while avoiding premature downward force, only allowing insertion once proper alignment is achieved. We also find that the residual policy makes subtle improvements to initial grasps, enabling more precise downstream alignment between the grasped object and the receptacle. Based on these observations, we posit that ResiP’s significant performance improvements stem from its ability to provide small adjustments in crucial moments—a natural fit for precision assembly tasks where minor misalignments often cause failure. While the base DP policy struggles with precise alignments and insertions, the learned residual component maintains proper alignment through minimal corrections.
However, the performance saturates at 70%-77% in the medium randomization settings, with parts initialized up to $\pm$5cm from nominal positions. This performance drop aligns with the intuition of the residual policy as improving precision and reactivity primarily as a local correction mechanism. When parts are placed far from their nominal positions, the DP policy may generate trajectories that deviate too far for the local correction mechanism to correct, leading to failures that are not easily corrected. Even if the failures can be corrected, the resulting states may be out-of-distribution for the DP policy, making corrections infeasible.
To better understand the failure modes of DP and ResiP in this higher randomness setting, we manually analyzed 75 randomly sampled trajectories for the round_table task on medium randomness (see the website for videos). We categorized the observed failures into four primary types:
  1. 1.
    ‘Peg Precision’ failures that occur during picking, inserting, or screwing operations with the table leg
  2. 2.
    ‘Foot Precision’ failures during picking, inserting, or screwing the table foot
  3. 3.
    ‘Out-of-Distribution’ failures when parts are unreachable by the base policy
  4. 4.
    ‘Other’ failures are attributed to contact-modeling issues in simulation (e.g., object penetration causing unrealistic accelerations) or timeout conditions.
Our qualitative analysis of these failure modes and recovery behaviors for the round_table task, shown in Fig. 7, shows that ResiP significantly improves reliability during precise insertion phases compared to the baseline approach, and eliminates the precision errors for “Peg Precision” entirely but that the ‘Out-of-Distribution’ increases in share.
Surprisingly, ResiP also leads to the emergence of qualitatively different behaviors compared to the nominal DP policy. We refer to the website’s uncut videos of rollouts for both policies. In these videos, we observe that ResiP has discovered alternative grasping strategies not present in the demonstration data—while demonstrations only showed grasping the table foot from above its center, ResiP learned to grasp from the side when the center was unreachable. Additionally, ResiP exhibits more decisive recovery behaviors, making deliberate adjustments that lead to successful task completion where the nominal BC policy would ineffectively oscillate back and forth.
Figure 8: Performance robustness when introducing random force perturbations during evaluation. At each timestep, we randomly apply forces to the objects in the environment, simulating unexpected disturbances. Our method (ResiP), which makes per-timestep corrections, shows greater resilience with only a 12% performance drop compared to nominal conditions. In contrast, chunk-based methods (DP, DP-DAgger, and ResiP-C) see larger drops of 19-26%. This difference in robustness highlights the value of closed-loop control for handling dynamic disturbances not seen during training. All methods were evaluated across 1024 episodes, both with and without perturbations.
IV-B What drives performance of ResiP?
@@ 40 Real+Sim 10/10 10/10 9/10 10/10 6/10 7/10 5/10 6/10 5/10 6/10
TABLE II: Comparing the effect of combining real-world demonstrations with simulation trajectories from our RL-trained residual policies. Co-training with real and synthetic data improves motion quality and success rate on the one_leg task.
@@
This section investigates different aspects of ResiP that improve task performance: (1) training stability and sample efficiency across RL methods, (2) the impact of addressing distribution shift through online data collection, (3) the benefits of closed-loop control compared to chunk-based execution, and (4) robustness to dynamic perturbations.
IV-B1 Performance and Training Characteristics of RL Methods
Our residual learning approach, ResiP, significantly improves upon baseline methods across all tasks. Using 50 demonstrations per task, where ResiP achieved 98% on the one_leg task (Tab. I), the alternative RL approaches show more limited improvement of 70% for PPO-C and 57% for IDQL. Similar trends hold for all tasks as seen in Tab. I.
We also see distinct training characteristics across methods. PPO-C, which directly fine-tunes the chunked MLP policy, exhibits training instability and requires careful KL regularization to avoid collapse [67, 68, 69]. IDQL faces a different challenge: even with maximum stochasticity in the denoising process, the base DP model produces actions with insufficient variance for effective exploration, limiting the potential for Q-learning-based improvement.
In contrast, ResiP demonstrates stable training behavior. Its architecture naturally constrains corrections to be local adjustments to the base policy’s absolute pose predictions rather than operating in the full workspace coordinate frame. This local prediction scope, combined with the residual network’s small parameter count, prevents the large policy updates that typically destabilize deep RL training [70, 71]. This stability provides consistent performance across hyperparameters (see Fig. 27 in App. J-C) and enables multiple gradient steps per collected trajectory, improving sample efficiency.
IV-B2 Impact of Distribution Shift
To probe the impact of mitigating distribution shift, we compared the baseline DP against DP-DAgger trained using online data collection via DAgger [72] as described in Sec. III-D. While DP-DAgger significantly outperforms the baseline DP (see yellow vs. red lines in LABEL:fig:overview (right) and Fig. 5), it still trails ResiP (blue line) by 9% and 23% for the one_leg and round_table tasks, respectively. The remaining performance gap suggests that reducing distribution shift with online data collection does not fully explain ResiP’s performance benefits.
IV-B3 Impact of Closed-Loop Control
To quantify the benefits of per-timestep corrections, we compared ResiP against ResiP-C, which predicts corrections at the chunk level as described in Sec. III-D4. Our experiments reveal that chunk-level corrections lead to significantly slower learning progress, with ResiP-C requiring significantly more environment interactions to achieve comparable performance (Fig. 25 in App. J-A). Moreover, even with extended training, ResiP-C’s performance saturates at 92% compared to ResiP’s 98% on the one_leg task. This performance gap suggests that the ability to make frequent corrections accelerates learning and enables higher terminal performance.
IV-B4 Robustness to Dynamic Perturbations
To further evaluate the benefits of closed-loop control, we compare the performance degradation of different methods under dynamic disturbances as described in Sec. III-C. As shown in Fig. 8, methods using chunk-based execution—DP, DP-DAgger, and ResiP-C—experience significant performance drops of 19-26 percentage points under these perturbations. In contrast, ResiP, with its ability to make per-timestep corrections, maintains a more robust performance with a 12% drop. This significant difference in robustness (7-14 percentage points) provides further evidence for the value of closed-loop control in precise manipulation tasks.
IV-C Real-World Deployment